# United Minds or Isolated Agents? Exploring Coordination of LLMs under Cognitive Load Theory

### v1

@@ B.4 Synthesizer Module: Consolidation and Cognitive Grounding
    • Cognitive Analogy: This setup mirrors Observational Learning (Bandura and Walters, 1977). The External Synthesizer observes the diverse problem-solving behaviors and refined outputs of the specialist agents. By analyzing these varied "models" of thought and their collective synthesis ($\mu^{(T-1)}$), it can construct a comprehensive solution, potentially integrating insights in a novel way without having been part of the iterative load distribution.
@@
This prompt structure ensures the Synthesizer has all necessary context to perform its role effectively.
Appendix C Experimental Setup: In-Depth Information
C.1 Detailed Benchmark Descriptions
LiveBench (White et al., 2025)
LiveBench serves as a dynamic and robust benchmark for evaluating LLM capabilities, characterized by its frequent updates (monthly) to minimize test data contamination and its focus on objectively scorable, challenging tasks. It draws from established hard benchmarks like Big-Bench Hard and AMPS, as well as introducing novel problems. The tasks span a broad range of domains, including:
  • Mathematics: Encompassing competitive programming problems, olympiad-level mathematics, and algebraic simplification.
  • Reasoning: Covering logical deduction and spatial reasoning.
  • Language: Focusing on nuanced understanding and manipulation.
  • Instruction Following: Testing adherence to complex instructions
  • Data Analysis: Requiring structured data manipulation
CommonGen-Hard (Madaan et al., 2023)
CommonGen-Hard, an extension of the CommonGen dataset (Lin et al., 2020), is specifically designed to impose high cognitive load by increasing element interactivity. The core task is to generate a coherent, multi-sentence paragraph incorporating a small set of 3-5 target concepts. The difficulty is amplified by including a large number (approximately 30) of irrelevant distractor concepts from which the model must select and use only the targets, while maintaining narrative coherence and commonsense plausibility. Given its generative nature, evaluation employs an LLM-based evaluator (gemini-1.5-pro) guided by a detailed rubric assessing ten dimensions. These dimensions are: (1) Relevance to Query (appropriateness and focus, highest weight); (2) Conciseness (brevity without losing essential content); (3) Clarity & Understandability (ease of comprehension); (4) Readability & Fluency (natural language flow, grammatical correctness); (5) Comprehensiveness & Completeness (addressing all prompt aspects); (6) Demonstrated Knowledge (accurate commonsense or domain knowledge); (7) Logic & Coherence (internal consistency and logical structure); (8) Originality & Creativity (novelty in ideas or framing); (9) Engagement & Interest (compelling nature of the response); (10) Insightfulness & Depth (analytical richness beyond surface content, lowest weight). Each dimension is scored (e.g., 1-10), and an aggregated total score is used. This setup directly tests the model’s ability to manage high element interactivity and filter relevant information, key aspects related to cognitive load.
C.2 Detailed Baseline Method Descriptions
The baseline methods used for comparison are implemented as follows:
@@ C.2 Detailed Baseline Method Descriptions
C.3 General Implementation Details
Experiments were conducted using Python and Google’s Generative AI SDK. LLM API Parameters: For all baseline methods (IO, CoT, SR) and the initial generation round ($t=0$) of multi-agent methods (MAD, DMAD, CoThinker), the API temperature was set to "0.25" to encourage some diversity. For subsequent iterative rounds ($t>0$) in CoThinker, MAD, and DMAD, the temperature was set to "0.0" and "frequency_penalty" to "0.5" to promote focused refinement and reduce repetition. Other API parameters (e.g., "top_p", "top_k") were left at their default values. Maximum output tokens were set appropriately for each task.
Appendix D Detailed Experimental Results and Ablation Studies
@@
D.1 Raw Subtask Performance Scores
@@ D.2 Subtask Descriptions
D.3 Ablation Study: Impact of Reference Set Size (N)
@@
D.4 Ablation Study: Impact of Exploration Rate (Beta)
@@
D.5 Ablation Study: Impact of Number of Agents (M)
@@
D.6 Ablation Study: Performance for Specific M/N
@@
Appendix E Limitations and Future Work
While CoThinker demonstrates promising results in managing cognitive load and enhancing collaborative LLM performance, this work has several limitations that also point towards avenues for future research.
Limitations include the scope of LLM evaluation, which primarily utilized models from the Gemini family. The generalizability of specific performance benefits and optimal hyperparameter settings across a wider range of LLM architectures requires further exploration. Additionally, while we argue that CoThinker manages transactional costs associated with multi-agent collaboration, a more fine-grained quantitative analysis of these costs versus the gains in solution quality would offer a more complete efficiency profile. The "thinking styles" currently rely on an LLM orchestrator and base styles; the true emergent specialization and their direct impact on distributing intrinsic load warrant deeper investigation.
Future Work could explore several promising directions. Developing adaptive CoThinker architectures that dynamically adjust parameters (number of agents, communication topology) based on real-time task assessment is a key area. Deeper integration of CLT principles, such as explicitly modeling and minimizing extraneous load from prompt design or fostering germane load via sophisticated scaffolding, could further enhance performance. Creating methods for explainability of collective cognition within CoThinker—tracing information flow, identifying critical contributions, and characterizing shared understanding evolution—would improve transparency. Extending the framework for human-AI collaboration, incorporating human users as specialized agents, could lead to powerful human-LLM group cognition. Finally, the prospect of such fused intelligence necessitates proactive examination of its societal implications, including equity, potential for misuse, accountability, and ethical considerations, demanding robust frameworks for responsible development and governance. Addressing these limitations and pursuing these future directions will further advance our understanding of how to build truly collaborative and cognitively capable LLM-based systems.

### v2

@@ D.6 Synthesizer Module: Consolidation and Cognitive Grounding
    • Cognitive Analogy: This setup mirrors Observational Learning (Bandura & Walters, 1977). The External Synthesizer observes the diverse problem-solving behaviors and refined outputs of the specialist agents. By analyzing these varied "models" of thought and their collective synthesis ($\mu^{(T-1)}$), it can construct a comprehensive solution, potentially integrating insights in a novel way without having been part of the iterative load distribution.
@@
This prompt structure ensures the Synthesizer has all the necessary context to perform its role effectively.
Appendix E Experimental Details and Results
Here we give detailed configuration of experiemnts from Section 5.1.
LLM API Parameters: For all baseline methods (IO, CoT, SR) and the initial generation round ($t=0$) of multi-agent methods (MAD, DMAD, CoThinker), the API temperature was set to "0.25" to encourage some diversity. For subsequent iterative rounds ($t>0$) in CoThinker, MAD, and DMAD, the temperature was set to "0.0" and "frequency_penalty" to "0.5" to promote focused refinement and reduce repetition. Other API parameters (e.g., "top_p", "top_k") were left at their default values. Maximum output tokens were set to be large enough for each task, otherwise specified in the task description.
All models run with the initial generation temperature set to 0.25 to encourage diverse outputs. In multi-agent settings, subsequent rounds use temperature 0.0 and a frequency penalty of 0.5 to reduce repetition. By default, multi-agent methods use M=6M{=}6 agents interacting over T=3T{=}3 rounds. For CoThinker, we set N=3N{=}3 references and exploration parameter $\beta{=}0.3$.
E.1 Detailed Baseline Method Descriptions
The baseline methods used for comparison from Section 5.1 are implemented as follows:
@@ E.1 Detailed Baseline Method Descriptions
Summary. (i) Single Agent (IO) is a standard mode of prompting without additional techniques. (ii) Single Agent (CoT) incorporates Chain-of-Thought prompting to elicit step-by-step reasoning. (iii) Single Agent (Self-Refine) uses iterative self-critique and revision processes. (iv) Multi-Agent Debate (MAD) employs interactive agent discussion with consensus formation. (v) Diverse MAD (DMAD) introduces heterogeneous prompting to avoid fixed mental sets and encourage diverse perspectives.
E.2 Benchmark Details
LiveBench (SOTA, comprehensive real-world tasks). LiveBench (White et al., 2025) is a widely used 2025 SOTA benchmark featuring challenging, objectively scorable tasks (e.g., high school math competitions, zebra and word puzzles) that demand high cognitive load for humans. It integrates content from established benchmarks including Big-Bench Hard (Suzgun et al., 2023), AMPS (Hendrycks et al., 2021b), and IFEval (Zhou et al., 2023). Frequent updates minimize test data contamination so results reflect genuine reasoning rather than memorization. Covered domains include:
  • Mathematics: competitive programming, olympiad-level math, algebraic simplification
  • Reasoning: logical deduction and spatial reasoning
  • Language: nuanced understanding and manipulation
  • Instruction Following: adherence to complex instructions
  • Data Analysis: structured data manipulation
The tasks are intentionally difficult—even strong models struggle—making LiveBench an ideal proxy for high cognitive load scenarios.
CommonGen-Hard (controlled information interactivity). CommonGen-Hard (Madaan et al., 2023), derived from CommonGen (Lin et al., 2020), systematically stresses element interactivity: the model must generate a coherent multi-sentence paragraph using 3–5 target concepts selected from a large pool (3̃0) of distractors while maintaining narrative coherence and commonsense plausibility. We adopt a 10-dimensional evaluation metric (Li et al., 2018) and employ an LLM-based evaluator (Gemini-1.5-Pro), following the protocol in Ye et al. (2023) to mitigate LLM-as-judge bias. The evaluator uses a detailed rubric with per-dimension Likert (1–10) scoring and aggregates to a total score. The ten rubric dimensions are:
  1. 1.
    Relevance to Query (appropriateness and focus; highest weight)
  2. 2.
    Conciseness (brevity without loss of essentials)
  3. 3.
    Clarity & Understandability (ease of comprehension)
  4. 4.
    Readability & Fluency (natural flow; grammar)
  5. 5.
    Comprehensiveness & Completeness (covers all prompt aspects)
  6. 6.
    Demonstrated Knowledge (accurate commonsense/domain knowledge)
  7. 7.
    Logic & Coherence (internal consistency and structure)
  8. 8.
    Originality & Creativity (novel framing/ideas)
  9. 9.
    Engagement & Interest (compelling responses)
  10. 10.
    Insightfulness & Depth (analytical richness; lowest weight)
This setup provides a controlled environment to validate that our architecture manages the information interactivity predicted by CLT to cause performance degradation, while keeping evaluation fair and systematic.
E.3 Cross-Model Detailed Results
This section provides comprehensive cross-model evaluation results referenced in Section 5.4, including detailed performance breakdowns across multiple LLM families and task categories.
Table 8 presents the full cross-model evaluation results across all tested models and evaluation scenarios. Table 9 provides detailed task-specific results for the standard setting cross-model evaluation, showing CoThinker’s performance across individual reasoning and mathematical tasks.
Model Method Avg. Score Math Score Reasoning Score
Gemini-2.5-Flash CoThinker 72.8 76.3 69.2
DMAD 59.7 56.7 62.8
IO (Baseline) 45.1 59.3 31.0
GPT-4.1-Mini CoThinker 55.4 40.0 70.8
DMAD 39.1 34.2 44.0
IO (Baseline) 37.4 34.0 40.8
Qwen3-32B CoThinker 22.1 18.9 25.2
DMAD 11.5 8.8 14.2
IO (Baseline) 11.7 3.4 20.0
DeepSeek-R1-8B CoThinker 5.8 2.9 8.8
DMAD 5.2 3.8 6.5
IO (Baseline) 2.3 1.9 2.8
Table 8: Complete cross-model evaluation results on LiveBench under constrained setting (8192 tokens, greedy decoding). CoThinker demonstrates consistent improvements across diverse model families and capabilities.
Model Configuration Zebra Puzzle Spatial Math Comp AMPS Hard
GPT-5-Nano (CoThinker) 75.75 88.00 89.13 88.00
GPT-5-Nano (IO) 56.75 80.00 78.26 87.00
Qwen3-30B-A3B (CoThinker) 83.50 84.00 84.78 88.00
Qwen3-30B-A3B (IO) 79.00 76.00 73.91 77.00
GPT-OSS-20B (CoThinker) 55.00 80.00 78.26 82.00
GPT-OSS-20B (IO) 54.00 78.00 71.74 82.00
Table 9: Task-specific performance breakdown across models and tasks. Full CoThinker configuration consistently outperforms baseline single-agent approach across reasoning and mathematical tasks, demonstrating the effectiveness of the complete multi-agent architecture.
The detailed results reveal several key insights about CoThinker’s generalization capabilities:
Model Family Independence: CoThinker demonstrates consistent improvements across architecturally diverse families (GPT, Gemini, Qwen, DeepSeek), indicating that the CLT-based approach generalizes beyond specific model architectures.
Performance Scaling: Stronger base models (e.g., GPT-5-Nano, Qwen3-30B-A3B) show larger absolute gains from CoThinker, suggesting that the framework effectively leverages higher baseline reasoning capabilities while providing substantial improvements for weaker models.
Task-Dependent Benefits: Mathematical reasoning tasks (math_comp, AMPS_Hard) and logical reasoning tasks (zebra_puzzle, spatial) consistently benefit from CoThinker’s collaborative approach, validating the cognitive load distribution hypothesis for high-intrinsic-load tasks.
Configuration Sensitivity: The comparison between full CoThinker configuration and baseline single-agent approach shows that the complete multi-agent architecture (including TMS, Communication Moderator, and Style Generation) provides meaningful improvements over simplified configurations.
E.4 Other Component Ablation Details
This section provides the complete component ablation results referenced in Section 5.6, showing detailed performance breakdowns for other model configurations tested. All configurations use M=6 agents and N=3 references with greedy decoding.
Task TMS: ON TMS: OFF
Math: Math Comp 35.6 7.7
Table 10: TMS ablation for Gemini-1.5-Flash-8B.
Task TMS: ON TMS: OFF
Math: Olympiad 47.1 44.2
Table 11: TMS ablation for Gemini-1.5-Flash.
Task TMS: ON TMS: OFF
Data Analysis: CTA 58.0 58.0
Instruction Following: Paraphrase 69.8 70.2
Instruction Following: Simplify 68.0 68.9
Instruction Following: Summarize 68.4 66.8
Language: Connections 53.5 52.6
Reasoning: Zebra Puzzle 48.2 39.5
Table 12: TMS ablation for Gemini-1.5-Pro.
Analysis: The TMS component shows variable effectiveness across model variants and task types. For the weaker Flash-8B model, TMS provides dramatic improvements on mathematical computation tasks, suggesting it helps manage cognitive load for complex reasoning. The standard Flash model demonstrates modest gains on advanced mathematical problems. The Pro model shows mixed results, with notable benefits for logical reasoning (Zebra Puzzle) but minimal impact on instruction-following tasks. This pattern indicates that TMS effectiveness depends on both model capability and task complexity, with greater benefits for cognitively demanding tasks on less capable models.
E.5 PPL Proxy Analysis for Component Ablation
This section (Table 13, Table 15, and Table 14) provides the complete set of PPL ablation studies referenced in Section 5.6.
Analysis.
  • Component effects (TMS and Style). Table 13 shows that both Thinking Style Orchestrator and TMS reduce PPL, indicating that each component lowers the cognitive effort required for the model to process peer information. Consistent with our captions, TMS yields the largest PPL relief, reflecting stronger scaffolding for integrating specialized knowledge.
  • Number of references ($k$). Table 14 shows PPL decreases as the number of peer answers increases, with diminishing returns beyond moderate $k$. This suggests that models find it easier to understand additional similar information, but overly large $k$ risks introducing unnecessary content reducing effective using of these answers, as explained in Section 4.3.
  • Selection strategy at k=3k{=}3. Table 15 shows that selecting similar peers achieves the lowest PPL, random also helps, while fully diverse peers increase PPL due to higher integration cost. As discussed in Section 4.3 (and Appendix D.5), prioritizing similarity reduces extraneous load but risks echo chambers; probabilistic rewiring ($\beta$) adds diversity to distribute intrinsic load while preserving small-world efficiency.
Model Base PPL +Style PPL +TMS PPL
Mistral-7B 6.6314 5.0422 1.5811
Qwen3-8B 6.5633 3.5814 1.6876
Table 13: PPL changes from baseline when adding Style or TMS. Both components reduce PPL; TMS shows the largest relief.
$k$ peers Mistral-7B PPL Qwen3-8B PPL
1 4.1516 3.3174
2 3.1343 3.1022
3 2.8578 3.0995
4 2.6123 2.5367
5 2.1221 2.1429
6 2.0917 1.8572
Table 14: PPL vs number of peer answers ($k$). Increasing $k$ consistently lowers PPL with diminishing returns beyond $k\geq 5$.
Selection (k=3k{=}3) Mistral-7B PPL Qwen3-8B PPL
Random (3rand) 2.0967 2.1924
Similar (3sim) 1.8891 1.5399
Diverse (3diverse) 3.0958 2.7905
Table 15: Effect of selection strategy at k=3k{=}3. Similar peers yield the lowest PPL; random also helps; fully diverse peers show higher PPL, consistent with higher integration cost.
Appendix F Detailed Experimental Results
@@
F.1 Raw Subtask Performance Scores
@@ F.2 Subtask Descriptions
F.3 Ablation Study: Impact of Reference Set Size (N)
@@
F.4 Ablation Study: Impact of Exploration Rate (Beta)
@@
F.5 Ablation Study: Impact of Number of Agents (M)
@@
F.6 Ablation Study: Performance for Specific M/N
@@
Appendix G Limitations and Future Work
Our work successfully demonstrates the utility of Cognitive Load Theory (CLT) as a generative framework for designing multi-agent LLM systems. The limitations of this initial study are best understood as defining the current boundaries of our methods and charting a course for developing more robust tools for this new area of research.
Developing Universal Proxies for Cognitive Load.
A central challenge is finding better tools to probe the internal cognitive states of LLMs. We introduced attention entropy and perplexity as initial, operational proxies for cognitive load, successfully showing their utility in guiding our architectural design. However, these measures are not yet a universal proxy; their values can be influenced by factors like model architecture and tokenization. A critical direction for future work is the search for more stable and universal measures, moving beyond static metrics to dynamic, process-based indicators that could form a standardized "cognitive toolkit" for quantifying the mental effort of artificial agents.
Boundary Conditions of the CLT Framework.
Our findings help delineate the boundary conditions under which our CLT-based model applies. The framework’s benefits are most pronounced for tasks with high intrinsic cognitive load, where the division of cognitive labor is essential. This suggests a compelling avenue for future work in developing a cognitive boundary conditions for LLMs, which would predict a priori which tasks require collaborative architectures.
Emergent Collective Dynamics.
While CoThinker operationalizes mechanisms for collective cognition, the rich, internal social dynamics remain a black box. A fascinating direction for future research is to apply methods from computational social science to analyze the interaction patterns within CoThinker, studying the evolution of the communication network or the emergence of transient leadership roles to better understand artificial collective intelligence.
Human-AI Collaboration.
A particularly exciting future direction is extending the framework to include human users as specialized agents. This could lead to powerful human-AI cognitive systems, where the architecture actively manages the cognitive load of both human and AI participants to solve problems that neither could tackle alone, fostering a true hybrid intelligence.
Societal and Ethical Implications.
Finally, the prospect of such powerful hybrid intelligence necessitates a proactive examination of its societal implications. Issues of equity, accountability, potential for misuse, and transparency demand robust ethical frameworks to ensure the responsible and beneficial development and governance of these systems.
Addressing these future directions will further advance our understanding of how to build truly collaborative and cognitively capable LLM-based systems.